find oxidation number for sulphur in H2so4 phosphorus in H#PO4 and PCI5​

Grade:11

1 Answers

Aastha Singh
3 years ago
Let X be the oxidation number of S inH2​SO4​.
The oxidation number of hydrogen is +1 and the oxidation number of oxygen is -2.
2(+1)+X+4(−2)=02+X−8=0X−6=0X=+6
Hence, the oxidation number of S inH2​SO4​is +6.
Similarly,
the oxidation numberphosphorus in H#PO4 is +7
the oxidation numberphosphorus inPCI5​ is +5